OpenStudy (anonymous):

any one know how to solve this problem? 13-5m=4-6m

OpenStudy (anonymous):

m=-9 I'm pretty sure, want to know why?

OpenStudy (anonymous):

I'll go through it anyway. You move all terms containing m to the left, all other terms to the right.

OpenStudy (anonymous):

Then add '6m' to each side of the equation. 13 + -5m + 6m = 4 + -6m + 6m

OpenStudy (anonymous):

Combine like terms: -5m + 6m = 1m 13 + 1m = 4 + -6m + 6m

OpenStudy (anonymous):

-6m + 6m = 0 13 + 1m = 4 + 0 13 + 1m = 4

OpenStudy (anonymous):

Then add '-13' to each side of the equation. 13 + -13 + 1m = 4 + -13

OpenStudy (anonymous):

Combine like terms again 13 + -13 = 0 0 + 1m = 4 + -13 1m = 4 + -13

OpenStudy (anonymous):

4 + -13 = -9 1m = -9

OpenStudy (anonymous):

Divide each side by '1'. m = -9 That's your answer! Hope this helped

OpenStudy (anonymous):

ok ty 

OpenStudy (anonymous):

m=-9

OpenStudy (anonymous):

You are solving for m here @Johnny621 is right !

OpenStudy (anonymous):